OpenStudy (zasharra):

Write a rational equation based upon the word problem you created. If Vanessa takes 2 hours to complete her assignment and it takes her friend Trish 4 hours to complete the same assignment how long would it take Vanessa and Trish to complete the assignment together?

OpenStudy (study312):

4 hours I think

OpenStudy (zasharra):

I have to write a rational equation for the problem @Study312

OpenStudy (zasharra):

@DangerousJesse do you know?

OpenStudy (study312):

hmm

OpenStudy (dangerousjesse):

Do you know what a rational equation is?

OpenStudy (zasharra):

no

OpenStudy (dangerousjesse):

A ration equation has one or more fractions as terms.

OpenStudy (zasharra):

\[\frac{ 1 }{ 2} + \frac{ 1 }{ 4} = \frac{ 1 }{ x}\] is this right?

OpenStudy (zasharra):

@DangerousJesse would this be the rational equation?

OpenStudy (zasharra):

@whpalmer4 @triciaal @Gabylovesyou @mathmate

OpenStudy (dangerousjesse):

One second.

OpenStudy (zasharra):

@Ghostgate

OpenStudy (zasharra):

@whpalmer @triciaal @Gabylovesyou @mathmate

OpenStudy (zasharra):

@whpalmer4